OpenStudy (vortish):

use the method of substitution to solve the system of linear equations 6a=3-b 6a=4-b multiple answers are given (a) there is one solution and the solution is (ordered pair) (b) there are intfantly many solutions (c) no solution

OpenStudy (anonymous):

it should be (a)

OpenStudy (vortish):

artix if the answer is a what would the ordered pair be

OpenStudy (anonymous):

from eqn 1, express b in terms of a, u get b=-2a, substitute that into eqn 2 and u get a=1 and b =-2

OpenStudy (mathstudent55):

Solve the first eq for be by first subtracting 3 from both sides: 6a - 3 = -b Now multiply both sides by -1: -6a + 3 = b Now rewrite as b = -6a + 3 Since you now know that b = -6a + 3, insert that value of b into b in the second equation: 6a = 4 - (-6a +3) and solve for a 6a = 4 + 6a -3 Collect like terms on right side: 6a = 1 + 6a Now subtract 6a from both sides: 6a - 6a = 1 + 6a - 6a and you get 0 = 1 Since this is a flase statement, that means there is no solution.